OpenStudy (anonymous):

Mark went to an electronic store, purchased 2 CDs and 6 DVDs, and spent $156. Antonio went to the same store, bought 1 CD and 9 DVDs, and spent $210. How much does each CD sell for at the store? $12 $15 $18 $22

OpenStudy (anonymous):

$12 $15 $18 $22

OpenStudy (anonymous):

$12

OpenStudy (anonymous):

It is 12. If you go: Mark - 12x2=24 156-24=132 132/6=22 So DvDs are $22 Antonio - 22x9= 198 198+12=210

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

if cd is x and dvd is y, 2x+6y=156 x+9y=210 2x+18y= 420 - 2x+6y=156 _____________________ 12y = 264 y = $22 x+9(22) = 210 x+198=210 x = $12

OpenStudy (anonymous):

thanks!